Rose, who is 16, is four times as old as her brother. How old will Rose be when she is half as old as her brother?
Found 2 solutions by henew, Alan3354:
Answer by henew(13)


Please read the question again and restate it. I think there is something missing.
But, based on your statement of the question.
Present Age of Ross = R =16
Present Age of Brother= B=16/4=4
When Rose is half as old as her brother, she will be 4/2= 2 years.

Answer by Alan3354(69443)

Rose, who is 16, is four times as old as her brother. How old will Rose be when she is half as old as her brother?
--------------
She will always be older, never 1/2 as old.